Out of 25 students, 14 had the chicken pox vaccine, 17 had the measles vaccine, and 7 had both vaccines, How many students had neither vaccines?
Found 2 solutions by ikleyn, richwmiller:
Answer by ikleyn(52775)

.
The number of those who were vaccinated and had at least one of the 2 vaccines is

14 + 17 - 7 = 24.


Hence, only one, 25 - 24 = 1 was not vaccinated.


Answer by richwmiller(17219)

Sometimes it is easier to view the diagram
7+7+10=24
25-24=1
Here is the Venn diagram for the problem